OpenStudy (anonymous):

Solve: |2t + 5| = 3 A. {1, 4} B. {-1, -4} C. {-1, 4} D. {1, -4}

OpenStudy (anonymous):

\(2t+5=\pm 3\)

OpenStudy (anonymous):

\[2t =3-5\\ 2t=-3-5\]

OpenStudy (anonymous):

oh ok. So would it be {-1, -4} @DanielM_113